True or False

1) The purpose of a battery in a circuit is to provide the charges.  

2) Two 3 Ohm resistors wired in series to a 12 V battery will use the same amount of power as a single 6 Ohm resistor wired to the same battery.  

3) 200 V across your body is probably more dangerous than 2 A flowing through your body.
Answer #1

1. False

Charge are already in the electric wire so battery need to just push the charge to flow one terminal to another.

2. True

Two three ohm resistance in series is total 6 ohm equivalent so power = v2/R =144/6= 24 w

And when 6 ohm resistor used

Power = 144/6 = 24 w

So power is same.

3. Correct

Current flow through voltage difference.if body has more resistant then it will less flow of current of a given potential difference.

Please upvote my answer. Thanks